OpenStudy (anonymous):

4. The quotient of 5/31 divided by 15/23, reduced to the lowest fraction, is A. 75/373. B. 93/23 or 4 1/23. C. 115/465. D. 23/93

OpenStudy (anonymous):

\[\huge \frac{5}{31}\div \frac{15}{23} \] \[\huge \rightarrow \frac{5}{31}\times \frac{23}{15} \] \[\huge \rightarrow \frac{1}{31}\times \frac{23}{3} = \frac{ 23}{ 3 \times 31}= \frac{ 23}{ 93}\]

OpenStudy (anonymous):

Hence option D is the correct answer